OpenStudy (anonymous):

Convert the decimal to a fraction. Be sure to show all of your work. 0.24

OpenStudy (nenadmatematika):

=24/100=6/25

OpenStudy (anonymous):

How can I show my work?

OpenStudy (anonymous):

well we know that there is 100% so u would take 0.24 x 100 which would = 24/100